News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 16 March 2001 9.00am EST

During the last week, Chandra completed the observing schedule as planned.

Real-time procedures were completed on 10 March to uplink the medium eclipse patch to configure the electrical power system for the period between eclipse seasons, and on 13 March to test a new version of software used to interface with the NASA communications network. A routine Ephemeris update was completed on 14 March.

Of note last week was the Chandra Cycle 3 Proposal deadline at midnight on March 15. Over 700 proposals were received requesting a total in excess of 150 million seconds, or approximately a factor of 8 oversubscription in observing time.

A press release was issued on March 13 describing results from over a million seconds of observation of the Chandra Deep Field South. The announcement describes an early universe "teeming" with black holes and has generated significant interest from the press. See http://chandra.harvard.edu/press/ for details.

The planned observations for the coming week are shown below and include a coordinated observation of GRS1915+105 with the RXTE mission on 24 March. Real-time procedures will be conducted on March 18 and 21 to insert and retract the LETG.

------------------------------------------
HZ43                    HRC-S/LETG   Mar 18
EK DRA                  HRC-S/LETG
Radiation Belt                       Mar 19
1ES1927+654             HRC-S/LETG   Mar 20
Epsilon Eri             HRC-S/LETG   Mar 21
Radiation Belt                       Mar 22
HE 2-10                 ACIS-S       Mar 23
NGC3226                 ACIS-S
3C303                   ACIS-S
1510-089                ACIS-S
AQL X-1                 ACIS-S
LBQS 1443+0141          ACIS-S
PC 1247+3406            ACIS-S
NGC4548                 ACIS-S
NGC4550                 ACIS-S
1RXSJ1229+1816          HRC-S        Mar 24
1RXSJ1452+3235          HRC-S
1RXSJ1328-3654          HRC-S
GRS 1758-258            HRC-S
GRS 1758-258            ACIS-S/HETG
GRS 1915+105            ACIS-S/HETG
Radiation Belt                       Mar 25
Hydra A                 ACIS-S/HETG

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports